# Estimation of Out-of-Sample Sharpe Ratio for High Dimensional Portfolio Optimization

### v1

@@ 6 Conclusion
In this work, we introduce a novel in-sample approach to estimate the out-of-sample Sharpe ratio in high-dimensional portfolio optimization, both with and without the risk-free asset, when the population covariance matrix ${\bm{\Sigma}}$ is unknown. By leveraging random matrix theory, we show that the proposed estimator shares the same asymptotic properties as the true Sharpe ratio, even when the operator norm of ${\bm{\Sigma}}$ is unbounded. Specifically, we only need the trace norm of ${\bm{\Sigma}}/p$ to be bounded, which makes our methodology more applicable to the real world financial data. We have also validated our theory through simulations and demonstrated its practical application through real data experiments.
In future investigations, it would be interesting to explore the convergence rate of the proposed estimator and determine if there are alternative estimators that offer faster convergence. This would involve studying the rate at which our estimator approaches the true Sharpe ratio as the sample size increases, and comparing it to other potential estimators. Understanding the convergence properties of different estimators can help identify optimal approaches for estimating the Sharpe ratio in different scenarios.
@@
Appendix A Alternative Settings for Numeric Experiments
In this section, we change the values of the population covariance matrix ${\bm{\Sigma}}$, the regularization matrix $\mathbf{Q}$ and the mean vector ${\bm{\mu}}$ separately to assess the accuracy of Theorem 2.7, while keeping the values of the other quantities the same as those in Section 4.1. This provides comparison as we evaluate the effects of ${\bm{\Sigma}}$, $\mathbf{Q}$ and ${\bm{\mu}}$ separately. We investigate these changes for $(n,p)=(1500,750)$ and $(n,p)=(1500,2250)$ as in the basic settings. We still generate $\mathbf{R}\in\mathbb{R}^{n\times p}$ for 100 times and follow the procedure in Section 4.1 to get $SR(\mathbf{Q})$ and $\widehat{SR}(\mathbf{Q})$.
@@ Appendix A Alternative Settings for Numeric Experiments
Results with different $\mathbf{Q}$’s. Here we fix ${\bm{\mu}}={\bm{\mu}}_{0}$ and ${\bm{\Sigma}}={\bm{\Sigma}}_{0}$. In the basic settings, we define $\mathbf{Q}=q\cdot\mathbf{Q}_{0}=q\cdot{\rm diag}(3,...,3,1,...,1)$, where the number of $3$’s and $1$’s are both $p/2$. To provide a comparison, we introduce two additional matrices, $\mathbf{Q}_{1}=0.5\mathbf{I}_{p}+q\cdot\mathbf{Q}_{0}$ and $\mathbf{Q}_{2}=2\mathbf{I}_{p}+q\cdot\mathbf{Q}_{0}$, when $(n,p)=(1500,750)$ thus $c=1/2$. We can treat $\mathbf{Q}_{1},\mathbf{Q}_{2}$ as selection of another parameter $\widetilde{q}\in\{0,0.5,2\}$ in $\mathbf{Q}=\widetilde{q}\mathbf{I}_{p}+q\mathbf{Q}_{0}$. Similarly, we consider $\mathbf{Q}_{1}=2\mathbf{I}_{p}+q\cdot\mathbf{Q}_{0}$ and $\mathbf{Q}_{2}=4\mathbf{I}_{p}+q\cdot\mathbf{Q}_{0}$ when $(n,p)=(1500,2250)$ thus $c=3/2$. The inclusion of the term $\mathbf{I}_{p}$ with different coefficients serves as a universal variance penalty, which makes the covariance matrix more homogeneous across diagonal entries. The simulation procedure is similar to Section 4.1, except that we now replace the regularization $q\cdot\mathbf{Q}_{0}$ with $\mathbf{Q}=\mathbf{Q}_{1}$ or $\mathbf{Q}=\mathbf{Q}_{2}$.
@@
In Figure 8, it can be observed that $\widehat{SR}(\mathbf{Q})$ matches with $SR(\mathbf{Q})$ for both $\mathbf{Q}=\mathbf{Q}_{1}$ and $\mathbf{Q}=\mathbf{Q}_{2}$ when $c=1/2$ and $c=3/2$. However, with different structures of $\mathbf{Q}$, the maximum value of SR can vary significantly. For instance, when $c=1/2$, the maximum value of SR in the basic settings is approximately $0.94$. However, in Figures 8(a) and 8(b), the maximum SR exceeds $1$. On the other hand, when $c=3/2$, the maximum value of SR in the basic settings is more than $0.97$. But in Figures 8(c) and 8(d), the maximum value of SR only approaches $0.92$. These results demonstrate the significance of determining the structure of the matrix $\mathbf{Q}$ in order to achieve a higher SR.
@@ B.1 Global minimum variance portfolio
In this section, we consider the global minimum variance portfolio, which has the advantage of no requirement about the knowledge of ${\bm{\mu}}$. Our construction of the global minimum variance portfolio is defined as $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}$, where we simply apply the regularization $\mathbf{Q}=q\mathbf{I}$. Since the total weights sum up to 1, the exact weight vector $\mathbf{w}$ takes the form:
$\mathbf{w}=\frac{(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}.$ (B.1)
Note that the scale is determined slightly differently from the mean variance portfolio, for the reason that ${\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}$ is always positive in the global minimum variance portfolio, but ${\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\mu}}$ may become negative and unstable in the mean-variance portfolio. Consequently, the variance of the portfolio can be expressed as:
$\operatorname{{\rm Var}}_{\widetilde{\mathbf{R}}}(\mathbf{w}^{\top}\widetilde{\mathbf{R}})=\frac{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}{({\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}})^{2}},$
where $\widetilde{\mathbf{R}}$ is an out-of-sample return vector with the distribution $\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$. Theorem 2.7 provides an in-sample method to estimate
$1/\sqrt{\operatorname{{\rm Var}}_{\widetilde{\mathbf{R}}}(\mathbf{w}^{\top}\mathbf{x})}=\frac{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}{\sqrt{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}},$
which exactly has the form of (2.1), replacing ${\bm{\mu}}$ with ${\bm{1}}$. To minimize the out-of-sample $\operatorname{{\rm Var}}_{\widetilde{\mathbf{R}}}(\mathbf{w}^{\top}\mathbf{x})$, Theorem 2.7 can help us determine the best $q$ from a finite candidate set. Below we give the steps of our real data analysis for the global minimum portfolio.
@@
    $\mathbf{Q}=q\mathbf{I}$. Consider two candidate sets $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$: $\mathcal{Q}_{1}=[1:30]/5$ for $c<1$ and $\mathcal{Q}_{1}=[1:30]/1.5$ for $c>1$; $\mathcal{Q}_{2}=[1:30]/10$ considers a narrower range of values, resulting in smaller gaps thus more refined differences among values of $q$.
@@
    For each testing month, we run experiments for all candidate $q$ values and also consider no regularization, i.e. $q=0$, where we have $\mathbf{w}\propto\widehat{\bm{\Sigma}}^{+}{\bm{1}}$ and $\widehat{\bm{\Sigma}}^{+}$ is the pseudo inverse, and the optimized $q^{*}\in\mathcal{Q}$ using the estimation in Theorem 2.7. Here,
    $q^{*}=\operatorname{argmax}_{q\in\mathcal{Q}}\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}\bigg{)}\cdot\frac{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}{\sqrt{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}}.$
@@
    We roll the procedure above for all testing months. Note that the value of $q^{*}$ changes from month to month. With the weight vector $\mathbf{w}$ using all $q\in\mathcal{Q}$, $q=0$ or $q^{*}$, we can then compute the portfolio returns for each trading day in the testing month.
  5. 5.
@@
    We report the standard deviation of the portfolio returns every three years, as shown in Figure 10. For instance, the results labeled on the x-axis as 202001 in Figure 10 represent the standard deviation of out-of-sample portfolio returns from January 2020 to December 2022 with 36 testing months. When the x-axis label updates to 202007, the results reflect the standard deviation of returns from July 2020 to June 2023.
@@
Figure 10: Standard deviation (volatility) of global minimum variance portfolios. The x-axis represents the rolling time, while the y-axis represents the standard deviation of the portfolio returns every three years. The blue solid line, orange dash-dot line and green dash line in the figure correspond to the standard deviation with $q=0$, $q=\min_{q\in\mathcal{Q}}q$ and $q=\max_{q\in\mathcal{Q}}q$ respectively. The boxplot displays the standard deviation for all $q\in\mathcal{Q}$, and the red triangle indicates the standard deviation under optimized $q=q^{*}\in\mathcal{Q}$.

### v3

@@ 6 Conclusion
In this work, we propose a novel in-sample method to estimate the out-of-sample Sharpe ratio in high-dimensional portfolio optimization with or without a risk-free asset, under an unknown population covariance matrix ${\bm{\Sigma}}$. By random matrix theory, we show that the proposed estimator shares the same asymptotic properties with the out-of-sample Sharpe ratio, even when $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is unbounded. Specifically, we only need the bounded $\|{\bm{\Sigma}}/p\|_{{\mathrm{tr}}}$, which is more applicable to the real financial data. We have also validated our theory through simulations and demonstrated its practical application through real data experiments.
In future work, it would be valuable to explore the asymptotic distribution of the proposed estimator and develop statistical inference, including studying the estimator’s correlation with the out-of-sample Sharpe ratio. A robust inference methodology would further enhance understanding and provide robust tools for evaluating the Sharpe ratio.
@@
Appendix A Additional Numerical Experiments
A.1 Alternative settings
In this section, we change the values of the population covariance matrix ${\bm{\Sigma}}$, the regularization matrix $\mathbf{Q}$ and the mean vector ${\bm{\mu}}$ separately to assess the accuracy of Theorem 2.7, while keeping the values of the other quantities the same as those in Section 4.1. This provides comparison as we evaluate the effects of ${\bm{\Sigma}}$, $\mathbf{Q}$ and ${\bm{\mu}}$ separately. We investigate these changes for $(n,p)=(1500,750)$ and $(n,p)=(1500,2250)$ as in the basic settings. We still generate $\mathbf{R}\in\mathbb{R}^{n\times p}$ for 1000 times and follow the procedure in Section 4.1 to get $SR(\mathbf{Q})$ and $\widehat{SR}(\mathbf{Q})$.
@@ A.1 Alternative settings
Results with different $\mathbf{Q}$’s. Here we fix ${\bm{\mu}}={\bm{\mu}}_{0}$ and ${\bm{\Sigma}}={\bm{\Sigma}}_{0}$. In the basic settings, we define $\mathbf{Q}=q\cdot\mathbf{Q}_{0}=q\cdot{\rm diag}(3,...,3,1,...,1)$, where the number of $3$’s and $1$’s are both $p/2$. To provide a comparison, we introduce three additional matrices, $\mathbf{Q}_{1}=0.1\mathbf{Q}_{0}+q\cdot{\rm diag}(\lambda_{1},...,\lambda_{p})$ where $\lambda_{i}$ is defined in Section 4.1; $\mathbf{Q}_{2}=0.5\mathbf{I}_{p}+q\mathbf{Q}_{0}$ and $\mathbf{Q}_{3}=q{\bm{\Sigma}}_{0}$. The simulation process is similar to that in Section 4.1, except that we now replace the regularization term $q\cdot\mathbf{Q}_{0}$ with $\mathbf{Q}=\mathbf{Q}_{1}$, $\mathbf{Q}=\mathbf{Q}_{2}$ or $\mathbf{Q}=\mathbf{Q}_{3}$. Additionally, we plot the line $SR_{\max}=\sqrt{{\bm{\mu}}_{0}^{\top}{\bm{\Sigma}}_{0}^{-1}{\bm{\mu}}_{0}}$ to compare the values of $SR(\mathbf{Q})$ and $SR_{\max}$, which provides further validation of the findings in Appendix F later and offers insight into how $\mathbf{Q}$ can be shaped to approximate the highest achievable Sharpe ratio.
@@
In Figure 8, it is clear that $\widehat{SR}(\mathbf{Q})$ aligns closely with $SR(\mathbf{Q})$ for both $\mathbf{Q}=\mathbf{Q}_{1}$ and $\mathbf{Q}=\mathbf{Q}_{2}$ when $c=1/2$ and $c=3/2$. However, the structure of $\mathbf{Q}$ significantly affects the maximally achievable Sharpe ratio. For instance, in the settings of $\mathbf{Q}=q\cdot\mathbf{Q}_{0}$ or $\mathbf{Q}=\mathbf{Q}_{2}$, the gap between the highest achievable Sharpe ratio $SR(\mathbf{Q})$ and the theoretically maximal Sharpe ratio $SR_{\max}$ is clearly noticeable. Interestingly, in the case of $\mathbf{Q}=\mathbf{Q}_{1}$ or $\mathbf{Q}=\mathbf{Q}_{3}$, as we adjust $\mathbf{Q}$ to approximate the structure of the population covariance matrix ${\bm{\Sigma}}$ via its residual covariance ${\rm diag}(\lambda_{1},...,\lambda_{p})$, or make $\mathbf{Q}$ exactly the same as ${\bm{\Sigma}}$, the maximal Sharpe ratio that we can achieve approaches $SR_{\max}$ as the tuning parameter $q$ increases. This observation not only supports our findings in Appendix F but also suggests a practical approach for designing $\mathbf{Q}$. Specifically, by constructing $\mathbf{Q}$ to be structurally similar or proportionally aligned with ${\bm{\Sigma}}$, we can enhance the Sharpe ratio and achieve values closer to $SR_{\max}$. Therefore, for optimal performance, it is advisable to design $\mathbf{Q}$ within the candidate set $\mathcal{Q}$ to reflect the true structure of ${\bm{\Sigma}}$ or its corresponding residual covariance. If we do not have prior information about the covariance structure, we may either use some independent data to estimate the residual covariance or find good proxies such as the trading volume to capture the rough level of the residual variances.
@@
A.2 Experiments with unknown mean vector
In this section, we present additional experiments on the situation where the mean vector ${\bm{\mu}}$ is unknown. Recall $\widehat{\bm{\Sigma}}$ and $\bm{\widehat{\mu}}$ are sample covariance and sample mean, respectively. The out-of-sample Sharpe ratio is given by $SR(\mathbf{Q})=\frac{\bm{\widehat{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{\bm{\widehat{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}}}$, and the estimation $\widehat{SR}(\mathbf{Q})$ is given by
$\widehat{SR}(\mathbf{Q})=\frac{\bm{\widehat{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}-\frac{{\mathrm{tr}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}}{n-{\mathrm{tr}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}}}{\sqrt{\bm{\widehat{\mu}}^{\top}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}}}\cdot\big{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\big{)}.$
See Appendix G for more details on the theoretical justification. We continue to fix the population matrix at ${\bm{\Sigma}}={\bm{\Sigma}}_{0}$ and conduct experiments across three scenarios with mean vectors ${\bm{\mu}}={\bm{\mu}}_{0}$, ${\bm{\mu}}_{1}$, and ${\bm{\mu}}_{2}$. For the regularization matrix $\mathbf{Q}$, we define it as $\mathbf{Q}_{1}=q\cdot{\rm diag}(\lambda_{1},\ldots,\lambda_{p})$ and $\mathbf{Q}_{2}=q\cdot{\bm{\Sigma}}_{0}$, where ${\bm{\Sigma}}_{0}$ and each $\lambda_{i}$ are specified in Section 4.1. Additionally, we plot the line corresponding to the maximal Sharpe ratio, given by $SR_{\max}=\sqrt{{\bm{\mu}}^{\top}{\bm{\Sigma}}^{-1}{\bm{\mu}}}$, to assess whether designing $\mathbf{Q}$ as approximately proportional to ${\bm{\Sigma}}$ can still lead to an approximation of $SR_{\max}$ when the mean vector ${\bm{\mu}}$ is unknown. Based on the discussion in Section G.2, we also plot the value $SR_{\text{L}}=\frac{SR_{\max}^{2}}{\sqrt{SR_{\max}^{2}+c}}$ where $SR(q\cdot{\bm{\Sigma}}_{0})$ approximate with large $q$. The constant $c$ here also reflects the high-dimensional nature of the problem: if $p$ is much smaller than the sample size $n$, it may be possible to approximate the maximal Sharpe ratio even when ${\bm{\mu}}$ is unknown.
Figure 10: Simulation results with ${\bm{\mu}}$ unknown. The x-axis and the y-axis follow exactly Figure 1.
The results presented in Figure 10 illustrate that when the mean vector ${\bm{\mu}}$ is unknown, our proposed estimator $\widehat{SR}(\mathbf{Q})$ closely aligns with the true Sharpe ratio $SR(\mathbf{Q})$. However, a particularly intriguing observation arises: in scenarios where ${\bm{\mu}}$ is known, a larger $q$ aligned with ${\bm{\Sigma}}$ seems beneficial for optimal portfolio construction; unlike the scenario where ${\bm{\mu}}$ is known, when we set $\mathbf{Q}=q{\bm{\Sigma}}$ in the case unknown ${\bm{\mu}}$, an increase in the tuning parameter $q$ maintains a gap between $SR_{\max}$ and the optimally achievable $SR(\mathbf{Q})$. Notably, the optimal value of $q$ may even appear in the intermediate range when $\mathbf{Q}$ has a slightly different form (change $\mathbf{Q}_{2}$ to $\mathbf{Q}_{1}$), suggesting that simply increasing $q$ does not always yield better results. This gap also indicates a challenge in approximating $SR_{\max}$ when using both sample mean and sample covariance together. In the context of an unknown ${\bm{\mu}}$, it appears that the best portfolio construction strategy may no longer involve using the true ${\bm{\Sigma}}$. As direct evidence, Figure 10(f) shows that with $\mathbf{Q}=q\mathbf{Q}_{1}$, where $\mathbf{Q}_{1}$ slightly differs from the true covariance matrix ${\bm{\Sigma}}$, the performance of $SR(\mathbf{Q}_{1})$ can even outperform the performance when $\mathbf{Q}=q\cdot{\bm{\Sigma}}$. This raises fascinating questions about the effectiveness of traditional approaches and suggests that when faced with uncertainty regarding ${\bm{\mu}}$, relying on the true covariance matrix might not lead to the best outcomes. Instead, an alternative approach may be necessary to navigate this uncertainty effectively. We leave this problem for future investigation.
A.3 Additional experiments on regularization matrix
In this section, we conduct additional experiments by optimizing the entire matrix $\widehat{\mathbf{Q}}=\arg\max_{\mathbf{Q}}\widehat{SR}(\mathbf{Q})$ (i.e. without restricting $\mathbf{Q}\in\mathcal{Q}$). The experiment is designed to support our claim below Theorem 2.7 that optimizing $\widehat{SR}(\mathbf{Q})$ over the entire space can lead to overfitting to the in-sample data, which will potentially inflate $\widehat{SR}(\widehat{\mathbf{Q}})$ and create a discrepancy between the estimated $\widehat{SR}(\widehat{\mathbf{Q}})$ and the true $SR(\widehat{\mathbf{Q}})$.
Recall that
$\widehat{SR}(\mathbf{Q})=\left(1-\frac{c}{p}\operatorname{tr}\left[\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\right]\right)\cdot\frac{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}}.$
To provide an example of overfitting where optimizing $\widehat{SR}(\mathbf{Q})$ without constraints, we specifically pick up two pairs of $(n,p)$: $(500,250)$ and $(1000,500)$ (other pairs are omitted due to heavy computational burden). We adopt the same settings as in Section 4.1 (${\bm{\mu}}={\bm{\mu}}_{0}$ and ${\bm{\Sigma}}={\bm{\Sigma}}_{0}$), generate $\mathbf{R}\in\mathbb{R}^{n\times p}$ for 20 times and optimize $\mathbf{Q}$ in $\widehat{SR}(\mathbf{Q})$ to obtain $\widehat{\mathbf{Q}}$ each time. To ensure that $\mathbf{Q}$ remains positive definite during optimization, we decompose $\mathbf{Q}$ as $\mathbf{Q}=\mathbf{L}\mathbf{L}^{\top}$ and then optimize $\mathbf{L}$, where $\mathbf{L}$ is a lower-triangular matrix. Additionally, we consider an alternative case by constraining $\mathbf{Q}$ to be diagonal.
Table 1: Comparison for $SR(\widehat{\mathbf{Q}})$ and $\widehat{SR}(\widehat{\mathbf{Q}})$. The mean gives the average value and the range gives the minimum and maximum values over the 20 independent trials.
$(n,p)$ $SR_{\max}$ Optimization over full $\mathbf{Q}$
mean of $SR(\widehat{\mathbf{Q}})$ Range of $SR(\widehat{\mathbf{Q}})$ mean of $\widehat{SR}(\widehat{\mathbf{Q}})$ Range of $\widehat{SR}(\widehat{\mathbf{Q}})$
$(500,250)$ 0.923 0.643 [0.604,0.694] 1.299 [1.178,1.454]
$(1000,500)$ 1.123 0.791 [0.738,0.824] 1.513 [1.406,1.578]
$(n,p)$ $SR_{\max}$ Optimization over diagonal $\mathbf{Q}$
mean of $SR(\widehat{\mathbf{Q}})$ Range of $SR(\widehat{\mathbf{Q}})$ mean of $\widehat{SR}(\widehat{\mathbf{Q}})$ Range of $\widehat{SR}(\widehat{\mathbf{Q}})$
$(500,250)$ 0.923 0.770 [0.715,0.818] 0.967 [0.909,1.056]
$(1000,500)$ 1.123 0.944 [0.912,0.979] 1.146 [1.081,1.224]
As shown in Table 1, when we optimize the full matrix $\mathbf{Q}$, the value of $\widehat{SR}(\widehat{\mathbf{Q}})$ is even much higher than $SR_{\max}$. Despite the optimal matrix $\widehat{\mathbf{Q}}$ having a high value in $\widehat{SR}(\widehat{\mathbf{Q}})$, the performance of the true objective $SR(\widehat{\mathbf{Q}})$ is poor, remaining significantly lower than $SR_{\max}$. This discrepancy indicates that the matrix $\widehat{\mathbf{Q}}$, which is optimized based on $\widehat{SR}(\mathbf{Q})$, is not suitable to obtain a high out-of-sample Sharpe ratio. As discussed below Theorem 2.7, the primary reason for discrepancy is that $\widehat{\mathbf{Q}}$ is overfitted to in-sample data. We further conduct experiments by restricting $\mathbf{Q}$ to be diagonal. Although this restriction reduces the discrepancy relative to the full matrix optimization, it does not eliminate it completely. With less flexibility in the optimization of $\mathbf{Q}$, the value of $\widehat{SR}({\mathbf{Q}})$ decreases, but the performance of $SR(\widehat{\mathbf{Q}})$ improves. In Appendix I, we prove that if $\mathbf{Q}$ is drawn from a fixed, finite‑dimensional family (fixed degree of freedom), then $\widehat{SR}(\widehat{\mathbf{Q}})$ and $SR(\widehat{\mathbf{Q}})$ are asymptotically consistent under mild conditions. It will be interesting to further investigate the relation between discrepancy and the flexibility of optimization in future work.
@@ B.1 Global minimum variance portfolio
In this section, we consider the global minimum variance portfolio, which has the advantage of no requirement about the knowledge of ${\bm{\mu}}$. Our construction of the global minimum variance portfolio is defined as $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}$, where we simply apply the regularization $\mathbf{Q}\in\mathcal{Q}_{1}$ or $\mathbf{Q}\in\mathcal{Q}_{2}$, where $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ are defined at beginning of Section 5. Since the total weights sum up to 1, the exact weight vector $\mathbf{w}$ takes the form:
$\mathbf{w}=\frac{(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}}{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}}.$ (B.1)
Note that the scale is determined slightly differently from the mean variance portfolio, for the reason that ${\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}$ is always positive in the global minimum variance portfolio, but ${\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}$ may become negative and unstable in the MV portfolio. Consequently, the variance of the portfolio can be expressed as:
$\operatorname{{\rm Var}}_{\widetilde{\mathbf{R}}}(\mathbf{w}^{\top}\widetilde{\mathbf{R}})=\frac{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}}{({\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}})^{2}},$
where $\widetilde{\mathbf{R}}$ is an out-of-sample return vector with mean ${\bm{\mu}}$ and covariance ${\bm{\Sigma}}$. Theorem 2.7 provides an in-sample method to estimate
$1/\sqrt{\operatorname{{\rm Var}}_{\widetilde{\mathbf{R}}}(\mathbf{w}^{\top}\mathbf{x})}=\frac{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}}{\sqrt{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}}},$
which exactly has the form of (2.1), replacing ${\bm{\mu}}$ with ${\bm{1}}$. To minimize the out-of-sample $\operatorname{{\rm Var}}_{\widetilde{\mathbf{R}}}(\mathbf{w}^{\top}\mathbf{x})$, Theorem 2.7 can help us determine the best $\mathbf{Q}^{*}$ from a finite candidate set. Below we give the steps of our real data analysis for the global minimum portfolio.
@@
    For each testing month, we run experiments for all candidate $\mathbf{Q}$ and also consider no regularization, i.e. $\mathbf{Q}={\bm{0}}$, where we have $\mathbf{w}\propto\widehat{\bm{\Sigma}}^{+}{\bm{1}}$ and $\widehat{\bm{\Sigma}}^{+}$ is the pseudo inverse, and the optimized $\mathbf{Q}^{*}\in\mathcal{Q}$ using the estimation in Theorem 2.7. Here,
    $\mathbf{Q}^{*}=\operatorname{argmax}_{\mathbf{Q}\in\mathcal{Q}}\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bigg{)}\cdot\frac{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}}{\sqrt{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}}}.$
@@
    We roll the procedure above for all testing months. Note that the value of $\mathbf{Q}^{*}$ changes from month to month. With the weight vector $\mathbf{w}$ using all $\mathbf{Q}\in\mathcal{Q}$, $\mathbf{Q}={\bm{0}}$ or $\mathbf{Q}^{*}$, we can then compute the portfolio returns for each trading day in the testing month.
@@
    We report the standard deviation of the portfolio returns every three years, as shown in Figure 11. For instance, the results labeled on the x-axis as 202001 in Figure 11 represent the standard deviation of out-of-sample portfolio returns from Jan 2020 to Dec 2022 with 36 testing months. When the x-axis label updates to 202007, the results reflect the standard deviation of returns from Jul 2020 to Jun 2023.
@@
Figure 11: Standard deviation (volatility) of global minimum variance portfolios. The x-axis represents the rolling time, while the y-axis represents the standard deviation of the portfolio returns every three years. The blue solid line, orange dash-dot line and green dash line in the figure correspond to the standard deviation with $q=0$, $q=0.1$ (minimum $q$) and $q=3$ (maximum $q$) respectively. The boxplot displays the standard deviation for all $\mathbf{Q}\in\mathcal{Q}$, and the red triangle indicates the standard deviation under optimized $\mathbf{Q}^{*}\in\mathcal{Q}$.